OpenStudy (anonymous):

Solve 4x - 2x - 2 = -42.

OpenStudy (anonymous):

x=20

OpenStudy (anonymous):

-20 i mean

OpenStudy (anonymous):

subtract ,like terms 4x-2x-2=-42 2x-2=-42 +2 +2 _________________ 2x= -40 \[\frac{2x}{2}=\frac{-40}{2}=-20\]

OpenStudy (anonymous):

4x-2x-2=-42 2x=-40 x=-40/2 x=-20